OpenStudy (anonymous):

[9.06] Create your own quadratic equation whose graph opens down and crosses the x-axis at two points. Use graphing technology to graph the parabola and approximate the coordinates of the vertex and x-intercepts. Provide your quadratic equation, the coordinates of the vertex, and the x-intercepts. Using complete sentences, explain the process of graphing the parabola using graphing technology and why using graphing technology is beneficial.

OpenStudy (anonymous):

Anybody reading this? :/

OpenStudy (anonymous):

have to open down or doesn't matter ?

OpenStudy (anonymous):

Must open down.

OpenStudy (anonymous):

\[y=-x ^{2}-10x-21\] vertex= -b/2a -(-10)/2(-1)=-5 substitution x=-5 find y \[y=-(-5)^{2}-10(-5)-21\]\[y=-25+50-21\] y=4 vertex= ( -5.4) set y=0 find x intercept \[-x ^{2}-10x-21 = 0\](x+7)(x+3)=0 x=-7 x=-3